# Two-sample comparison through additive tree models for density ratios

### v1

@@ 2.3 Choice of the function class: additive tree models
Under this model, we now are presented with the following optimization problem:
@@
Next, we introduce two algorithms for fitting the above model under the balancing loss. These are direct counterparts of two different approaches to fitting additive trees in supervised learning problems. The first corresponds to a forward-stagewise (FS) algorithm based on greedy steepest descent on the space of single tree-based piecewise-constant functions; the other corresponds to Friedman (2001)’s gradient boosting (GB) algorithm. In practice, these algorithms typically produce estimates that are very similar to each other, and so one could just adopt the simpler GB algorithm, but because the details of the FS algorithm shed more light on what constitutes effective function approximators for the density ratio under the balancing loss, we believe they are worthy of elaboration here.
@@
The following proposition characterizes the optimal single tree fit in each iteration of the boosting algorithm.
@@
Given the current estimate $w_{k-1}$, the loss is minimized when the following conditions are satisfied:
@@
  2. 2.
    Then, $T_{k}$ is a partition structure that maximizes the Hellinger distance
The proof is provided in Supplementary Materials A.
Remark 1: If the balancing weight is correctly estimated, it achieves the symmetry between the two expectations in the loss in that $\mathbb{E}_{p}[w^{-1}]=\mathbb{E}_{q}[w]$, and we can impose this condition as a restriction on the estimator so that it does not focus on minimizing only one expectation in the two. As its justification, Proposition 2.1 implies that the balancing loss can be improved by multiplying the current estimate $w_{k}$ by a constant ($c$, say) such that $\mathbb{E}_{p}[(cw_{k})^{-1}]=\mathbb{E}_{q}[cw_{k}]$ (see Supplementary Materials A for the detail). Based on the discussion, we introduce this correction step in the boosting algorithms.
Remark 2: Given the connection of the balancing loss to the variational form of the Hellinger divergence, it is not surprising that the “optimal” single-tree in each iteration corresponds to the partition that maximizes the corresponding discretized approximation to the Hellinger divergence.
The optimal learner in the popular gradient-based boosting is obtained based on the first order condition in Friedman et al. (2000), while the proposition gives a solution that exactly maximizes the reduction in the balancing loss, or equivalently the increase in Hellinger distance. This is also equivalent to minimizing the discretized affinity, $\sum_{A\in T_{k}}\sqrt{P_{k}|_{T_{k}}(A)Q_{k}|_{T_{k}}(A)}.$
While in practice searching for the exact optimal tree in each step is not computationally tractable, this inspires a greedy coarse-to-fine algorithm in which we iteratively choose a single cut point that achieves the steepest ascent in Hellinger distance or the steepest ascent in the corresponding affinity. Specifically, when we consider splitting the current node $A$, we can search for the splitting rule such that for the children nodes $A_{l}$ and $A_{r}$ the following affinity is minimized:
@@
We can also introduce a gradient-based tree learner (Friedman, 2001). If a current estimate is $w$, with the log-transformation
$F=\log w,$
we formulate an ensemble where learners are combined additively. Then we can construct a tree partition that fits the functional gradients (i.e., residuals) defined for the two groups. As the two integrands are written as $e^{-F(x)}$ and $e^{F(x)}$, the negative gradients are as follows:
$\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(0)})}=-\frac{1}{n_{0}}e^{-F(x_{i}^{(0)})}=\frac{1}{n_{0}}w^{-1}(x_{i}^{(0)})\quad\text{and}\quad-\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(1)})}=\frac{1}{n_{1}}e^{F(x_{i}^{(1)})}=-\frac{1}{n_{1}}w(x_{i}^{(1)}).$
We follow the recipe provided in (Friedman, 2001): we first fit the new tree partition to the gradients by minimizing squared error loss, that is, minimizing variances of the gradients in the leaf nodes. To this end, we adopt the coarse-to-fine greedy search and for each node $A$, we select the splitting rule with children nodes $A_{l}$ and $A_{r}$ that minimizes the variances. After constructing the tree, secondly, we obtain the optimal values of the node parameters, which are already given by Proposition 2.1.
2.4 Regularization strategies
The success of tree boosting under the two algorithms above requires each base learner to be weak in the sense that it extracts only a small amount of structure in the density ratio function. Similar to supervised boosting, we can regularize the contribution of each learner to the estimate by introducing a learning rate denoted by $\nu\in(0,1]$. If the current estimate is $w_{k-1}$ and the optimal learner obtained in the $k$th step is $f_{k}$, we can combine them so that
@@ 2.4 Regularization strategies
An additional, complementary strategy for regularization is to introduce a restriction on the size of trees to avoid constructing overly complicated partition structures, so we set the maximum resolution (depth) of the trees to be, for example, between 4 and 6. Finally, the number of trees in the ensemble also needs be carefully selected. We find that choosing this number via cross-validation yields satisfactory performance.
@@ 3.1 Balancing loss as a pseudo-likelihood
To achieve uncertainty quantification on the density ratio, we use the balancing loss as a pseudo-likelihood within the framework of the generalized Bayesian inference (Bissiri et al., 2016). Specifically, we define a pseudo-likelihood using the finite-sample loss $l_{n}(w)$, that is
@@
Next we describe prior specification on $w$ for efficient MCMC and a data-driven approach for setting the value of $\tau$.
@@ 3.2 Conjugate prior specification
Recall our model given in (4), we describe a Gibbs sampling strategy that iteratively samples the full conditional of each basis function $f_{k}$, given all the others. That is, how one may sample the node parameters $\beta_{k}(A)$ and the tree partition $T_{k}$ in that defines $f_{k}$ as in (5).
@@ 3.2.1 Updating leaf node parameters
To derive the full conditional for $\beta_{k}(A)$, we define the balancing weight from which $f_{k}$ is subtracted:
@@
The portion in the likelihood $L_{n,\tau}(w)$ that depends on $\beta_{k}(A)$ is given as
@@
We set $\mu_{k}(A)$ to $1$ so that the induced marginal prior on $\log w$ can be around 0 (more on this below). The precision parameter $\lambda_{k}(A)$ induces a shrinkage toward the prior mean 1, as seen in the posterior mean $\mu^{\prime}_{k}(A)$ given in Eq (6). We set $\lambda_{k}(A)$ according to the number of trees $K$:
@@
where $\lambda_{0}$ is a hyperparameter. The prior value of $\lambda_{k}(A)$ is linear in $K$ because the results provided in Whitmore and Yalovsky (1978) imply that the variance of $f_{k}(x)$ for each point $x$ ($x\in\Omega$) is $O(K^{-1})$ under our model specification (see Supplementary Materials A for more detail). This choice of $\lambda_{k}(A)$ is also analogous to the amount of shrinkage recommended for the BART model (Chipman et al., 2010). We set $\lambda_{0}$ to 5 as a default value in all of our numerical experiments.
Another consideration is that the inverse Gaussian distribution has an asymmetric distribution structure, so it can also introduce skewness to the prior of the log-balancing weight $\log w$. To avoid this issue, we introduce the inverse Gaussian prior to $\gamma_{k}(A)$ and $\gamma^{-1}_{k,A}$ alternately, that is,
@@
This also makes the marginal prior of $\log w$ exactly centered at 0 whenever the total number of trees $K$ is even.
3.2.2 Updating tree structures
@@
We set the hyperparameters to the widely used default choices as $a_{T}=0.95$ and $b_{T}=2$ (Chipman et al., 2010). When splitting the node, one of the $d$ candidate dimensions of the sample space is selected with uniform probabilities, and a cut point is selected uniformly over the range in that dimension. Following the standard recipe for BART (Chipman et al., 2010; Kapelner and Bleich, 2016; Sparapani et al., 2021), we update the tree $T_{k}$ by Metropolis-Hastings moves. The details on the tree updating moves are provided in Supplementary Materials B.
@@ 3.3 Tuning the temperature $\tau$
In generalized Bayesian inference, the temperature parameter $\tau$ controls the strength of the loss-based likelihood for the balancing weight function $w$ as opposed to the prior and significantly influences the resulting posteriors, such as posterior variances and credible intervals. There is a growing number of discussions on strategies for selecting the temperature parameter empirically. Some proposed methods include information matching (Holmes and Walker, 2017), calibrating frequentist coverage rates of credible intervals (Syring and Martin, 2019), and the SafeBayes algorithm (Grünwald and van Ommen, 2017). It is not easy to adapt these methods in the current context where the task is unsupervised with an unknown function of interest specified nonparametrically.
Instead, we adopt a hierarchical Bayesian approach by placing a prior on the temperature $\tau$, following Bissiri et al. (2016); Rigon et al. (2023). The full conditional of $\tau$ is proportional to
@@
where $\pi_{0}(\tau)$ corresponds to the prior for $\tau$. We adopt a conjugate prior $\pi_{0}(\tau)$, namely $\mathrm{Gamma}(a_{0,\tau},b_{0,\tau})$, so the full conditional of $\tau$ is $\mathrm{Gamma}(a_{1,\tau},b_{1,\tau})$, where
@@
The analytic form of the full conditional elucidates the role $\tau$ plays and justifies the data-dependent approach to tuning of its value. Specifically, $\tau$ can be interpreted as a regularization parameter that penalizes poor overlap of the two samples. To see this, note that the full conditional mean of $\tau$ is $\frac{a_{0,\tau}+n}{b_{0,\tau}+nl_{n}(w)}$, so when the total sample size is large, the inverse of this quantity approximates the realized value of the loss $l_{n}(w)$. When the two samples have little overlap, for example, the estimated weights $w^{-1}$ for the first group and $w$ for the second group tend to be both small and close to zero, leading to a small $l_{n}(w)$ and thus a large $\tau$. Conversely, when the samples have good overlap, the weights are closer to 1, yielding a larger $l_{n}(w)$ and a smaller $\tau$. As an illustration, Figure 1 shows the posterior distributions of $\tau$ under the different levels of overlap under a simple simulation experiment. The observations ($n_{0}=n_{1}=200$) are generated from the standard normal distribution, and the second group is mean-shifted by $1,2$, and $3$, corresponding to decreasing levels of overlap. We can see that poorer overlap of the two samples leads to larger absolute values of $\log w$, along with larger values of $\tau$ corresponding to the smaller realized value of the loss $l_{n}(w)$.
Figure 1: Posterior distributions of $\tau$ under different levels of overlap. First row: The black and red points correspond to the two samples, and the dashed lines indicate the posterior means of the log-density ratios. Second row: The density functions of the posterior distributions of $\tau$ are plotted.
@@ 4 Numerical Experiments
We conduct simulation experiments to evaluate the proposed methods—both tree boosting and Bayesian additive trees. We begin with scenarios involving bivariate distributions, and then consider higher-dimensional (20-dimensional) cases. We provide a comparison with several existing approaches. The first approach to be compared with is the “density-ratio trick” by inverting a binary classifier. In particular, we adopt AdaBoost as binary classification for it uses the same class of an additive tree ensemble to classify the two classes using essentially the exponential loss (Friedman et al., 2000). The difference between this approach and our boosting algorithms will essentially be due to the different loss function we adopt. To apply the density-ratio trick, we first estimate the classification probabilities and then transform them into the density ratios by multiplying the sample size ratios. For implementation, we use the ada function provided in the R package ada (Culp et al., 2016). In addition, besides the “density-ratio trick”, we also compare our methods with two popular kernel-based density ratio estimators, namely KLEIP (Sugiyama et al., 2008) and uLSIF (Kanamori et al., 2009), implemented with the R package densratio (Makiyama, 2019).
For boosting, we assess both the FS and the GB algorithms. The learning rate $\nu$ is set to 0.01, and the maximum resolution of each tree is set to 4. The optimal number of trees is selected with the 5-fold cross-validation with the maximum number allowed 1000. For a fair comparison, AdaBoost was implemented under the same tuning parameters. For the generalized Bayesian additive tree model, we set the size of the ensemble to 200, following the standard implementation of the BART (Chipman et al., 2010). The MCMC is initialized with trees with only roots, and the size of the burn-in and the posterior sample are set to 2000 and 1000, respectively. In both types of the proposed models, we introduce 31 cut points with equal gaps in each dimension. Further details on the implementation are provided in Supplementary Materials C.
For each scenario, we evaluate the performance under two sample size settings: (i) $n_{0}=n_{1}=5000$ (“balanced” setting) and (ii) $n_{0}=9000$, $n_{1}=1000$ (“unbalanced” setting). As a measure of the accuracy, we use the following symmetrized squared error for the true density ratio $r^{*}$ and the estimated ratio $\hat{r}$:
$\frac{1}{n_{0}}\sum^{n_{0}}_{i=1}\left(\log r^{*}(x_{i}^{(0)})-\log\hat{r}(x_{i}^{(0)})\right)^{2}+\frac{1}{n_{1}}\sum^{n_{1}}_{i=1}\left(\log r^{*}(x_{i}^{(1)})-\log\hat{r}(x_{i}^{(1)})\right)^{2},$
and for each setting we calculate the MSE based on 50 simulated data sets.
@@
We simulate data sets using the following three scenarios with various types of differences in the distributions of $X_{0}$ and $X_{1}$:

### v2

@@ 2.3 Additive tree models and boosting
Under this model, we are now presented with the following optimization problem:
@@
Next, we introduce two algorithms for fitting the above model under the balancing loss. These are direct counterparts of two different approaches to fitting additive trees in supervised learning problems. The first corresponds to a forward-stagewise (FS) algorithm based on greedy steepest descent on the space of single-tree-based piecewise-constant functions; the other corresponds to Friedman (2001)’s gradient boosting (GB) algorithm. In practice, these algorithms typically produce estimates that are very similar to each other, and so one could just adopt the simpler GB algorithm, but because the details of the FS algorithm shed more light on what constitutes effective function approximators for the density ratio under the balancing loss, we believe they are worthy of elaboration here.
@@
The following proposition characterizes the optimal single tree fit in each iteration of a forward-stagewise algorithm, which takes $K$ steps of iterative fitting and in each step $k$ finds the optimal fit of $f_{k}$ that minimizes the balancing loss. Its proof is provided in Supplementary Materials A.
@@
Suppose $w_{k-1}$ is the estimate of $w$ after (k1)(k-1) steps of fitting. Then in the $k$th step, the loss is minimized by $\log w_{k}=\log w_{k-1}+f_{k}$ when the following conditions are satisfied by $f_{k}$:
@@
    Then $T_{k}$ is a partition structure that maximizes the Hellinger distance
  2. 2.
@@
Remark 1: If the balancing weight is correctly estimated, it achieves the symmetry between the two expectations in the loss in that $\mathbb{E}_{p}[w^{-1}]=\mathbb{E}_{q}[w]$, and we can impose this condition as a restriction or regularization on the estimator so that it does not focus on minimizing only one expectation out of the two. As justification, Proposition 2.1 implies that the balancing loss can be improved by multiplying the current estimate $w_{k}$ by a constant ($c$, say) such that $\mathbb{E}_{p}[(cw_{k})^{-1}]=\mathbb{E}_{q}[cw_{k}]$ (see Supplementary Materials A for details). Based on this discussion, we introduce this correction step in the boosting algorithms.
Remark 2: Given the connection between the balancing loss and the variational form of the Hellinger divergence, it is not surprising that the “optimal” single-tree in each iteration corresponds to the partition that maximizes the corresponding discretized approximation of the Hellinger divergence.
As in tree-based predictive models, searching for the exact optimal tree in each step is often not computationally tractable for multivariate data. Instead one can adopt a greedy, coarse-to-fine algorithm in which we iteratively choose a single cut point at each split of the tree $T_{k}$ that achieves the steepest ascent in the Hellinger distance or the steepest ascent in the corresponding affinity. Specifically, when we consider splitting the current node $A$, we can search for the splitting rule such that for the child nodes $A_{l}$ and $A_{r}$ the following affinity is minimized:
@@
We can also introduce a gradient boosting algorithm (Friedman, 2001). To this end, note that the negative gradients of the empirical loss with respect to the additive estimand $F$ are as follows:
$\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(0)})}=-\frac{1}{n_{0}}e^{-F(x_{i}^{(0)})}=\frac{1}{n_{0}}w^{-1}(x_{i}^{(0)})\quad\text{and}\quad-\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(1)})}=\frac{1}{n_{1}}e^{F(x_{i}^{(1)})}=-\frac{1}{n_{1}}w(x_{i}^{(1)}),$
which are the pseudo-residuals. One can then follow the recipe provided in (Friedman, 2001). Specifically, in each iteration of the algorithm we fit a regression tree to the pseudo-residuals by minimizing the squared error loss, that is, minimizing the variances of the gradients in the leaf nodes. We again adopt a coarse-to-fine greedy search and for each node $A$, we select the splitting rule with child nodes $A_{l}$ and $A_{r}$ that minimizes the variances. After constructing the tree, we obtain the optimal values of the node parameters, which are already given by Proposition 2.1.
Regularization strategies
As for supervised problems, the success of tree boosting under the two algorithms above requires each base learner to be weak in the sense that it extracts only a small amount of structure in the density ratio function. Similar to supervised boosting, we can regularize the contribution of each learner to the estimate by introducing a learning rate denoted by $\nu\in(0,1]$. If the current estimate is $w_{k-1}$ and the optimal learner obtained in the $k$th step is $f_{k}$, we can combine them as follows:
@@
An additional, complementary strategy for regularization is to introduce a restriction on the size of trees to avoid constructing overly complicated partition structures, so we set the maximum resolution (depth) of the trees to be, for example, between 4 and 6. Finally, the number of trees in the ensemble also needs to be carefully selected. We find that choosing this number via cross-validation yields satisfactory performance.
@@ 3.1 Balancing loss as a pseudo-likelihood
Tree boosting only offers point estimates for the density ratio function, without providing any uncertainty associated. To achieve uncertainty quantification for the density ratio, we can instead adopt a (generalized) Bayesian strategy (Bissiri et al., 2016) to learning the density ratio. Specifically, we define a pseudo-likelihood using the finite-sample loss $l_{n}(w)$, that is
@@ 3.2 Conjugate prior specification
Next we describe prior specification on $w$ for efficient MCMC and a data-driven approach for setting the value of $\tau$. Recalling our model introduced in (4), we describe a Gibbs sampling strategy that iteratively samples the full conditional of each basis function $f_{k}$, given all the others. That is, how one may sample the node parameters $\beta_{k}(A)$ and the tree partition $T_{k}$ that define $f_{k}$ as in (5).
@@ 3.2.1 Updating leaf node parameters
To derive the full conditional for $\beta_{k}(A)$, we define the balancing weight with $f_{k}$ subtracted:
@@
The portion of the likelihood $L_{n,\tau}(w)$ that depends on $\beta_{k}(A)$ is given by
@@
We set $\mu_{k}(A)$ to $1$ so that the induced marginal prior on $\log w$ is around 0 (more discussions are to be given at the end of this section). The precision parameter $\lambda_{k}(A)$ induces shrinkage toward the prior mean 1, as seen in the posterior mean $\mu^{\prime}_{k}(A)$ given in Eq (6). We set $\lambda_{k}(A)$ according to the number of trees $K$:
@@
where $\lambda_{0}$ is a hyperparameter. The prior value of $\lambda_{k}(A)$ is linear in $K$ because the results in Whitmore and Yalovsky (1978) imply that the variance of $f_{k}(x)$ for each point $x$ ($x\in\Omega$) is $O(K^{-1})$ under our model specification (see Supplementary Materials A for more details). This choice of $\lambda_{k}(A)$ is also analogous to the amount of shrinkage recommended for the BART model (Chipman et al., 2010). We set $\lambda_{0}$ to 5 as a default value in all of our numerical experiments.
Another consideration is that the inverse Gaussian distribution has an asymmetric shape, so it can also introduce skewness into the prior of the log-balancing weight $\log w$. To avoid this issue, we apply the inverse Gaussian prior to $\gamma_{k}(A)$ and $\gamma^{-1}_{k,A}$ alternately, that is,
@@
This ensures that the marginal prior of $\log w$ is exactly centered at 0 whenever the total number of trees $K$ is even.
3.2.2 Tree priors and posterior updates on the trees
@@
We set the hyperparameters to the widely used default choices as $a_{T}=0.95$ and $b_{T}=2$ (Chipman et al., 2010). When splitting the node, one of the $d$ candidate dimensions of the sample space is selected with equal probability, and a cut point is selected uniformly over the range in that dimension. Following the standard recipe for BART (Chipman et al., 2010; Kapelner and Bleich, 2016; Sparapani et al., 2021), we update the tree $T_{k}$ by using Metropolis-Hastings moves. The details of the tree updating moves are provided in Supplementary Materials B.
@@ 3.3 Tuning the temperature $\tau$
In generalized Bayesian inference, the temperature parameter $\tau$ controls the strength of the loss-based likelihood for the balancing weight function $w$ relative to the prior and significantly influences the resulting posteriors, including posterior variances and credible intervals. There is a growing body of literature on strategies for selecting the temperature parameter empirically. Some proposed methods include information matching (Holmes and Walker, 2017), calibrating frequentist coverage rates of credible intervals (Syring and Martin, 2019), and the SafeBayes algorithm (Grünwald and van Ommen, 2017). However, it is not straightforward to adapt these methods to the current context, where the task is unsupervised and the function of interest is an unknown, nonparametrically specified function.
Instead, we adopt a hierarchical Bayesian approach by placing a prior on the temperature $\tau$, following Bissiri et al. (2016); Rigon et al. (2023). To this end, we add $\tau^{n}$ to the pseudo-likelihood so that the full conditional of $\tau$ is proportional to
@@
where $\pi_{0}(\tau)$ denotes the prior for $\tau$. We adopt a conjugate prior $\pi_{0}(\tau)$, namely $\mathrm{Gamma}(a_{0,\tau},b_{0,\tau})$, so the full conditional of $\tau$ is $\mathrm{Gamma}(a_{1,\tau},b_{1,\tau})$, where
@@
The analytic form of the full conditional elucidates the role $\tau$ plays and justifies the data-dependent approach to tuning its value. Specifically, $\tau$ can be interpreted as a regularization parameter that penalizes poor overlap of the two samples. To see this, note that the full conditional mean of $\tau$ is $\frac{a_{0,\tau}+n}{b_{0,\tau}+nl_{n}(w)}$, so when the total sample size is large, the inverse of this quantity approximates the realized value of the loss $l_{n}(w)$. When the two samples have little overlap, for example, the estimated weights $w^{-1}$ for the first group and $w$ for the second group tend to be both small and close to zero, leading to a small $l_{n}(w)$ and thus a large $\tau$. Conversely, when the samples have good overlap, the weights are closer to 1, yielding a larger $l_{n}(w)$ and a smaller $\tau$. As an illustration, Figure 1 shows the posterior distributions of $\tau$ under different levels of overlap under a simple simulation experiment. The observations ($n_{0}=n_{1}=200$) are generated from the standard normal distribution, and the second group is mean-shifted by $1,2$, and $3$, corresponding to decreasing levels of overlap. We can see that poorer overlap of the two samples leads to larger absolute values of $\log w$, along with larger values of $\tau$ corresponding to the smaller realized value of the loss $l_{n}(w)$.
Figure 1: Posterior distributions of $\tau$ under different levels of overlap. Top row: The black and red points correspond to the two samples, and the dashed lines indicate the posterior means of the log-density ratios. Bottom row: The density functions of the posterior distributions of $\tau$ are shown.
@@ 4 Numerical Experiments
We conduct simulation experiments to evaluate the proposed methods—both tree boosting and Bayesian additive trees. We begin with scenarios involving bivariate distributions, and then consider higher-dimensional (20-dimensional) cases. We provide a comparison with several existing approaches. The first approach to be compared with is the “density-ratio trick” by inverting a binary classifier. In particular, we adopt AdaBoost as binary classification, as it uses the same class of additive tree ensembles to classify the two classes using essentially the exponential loss (Friedman et al., 2000). The difference between this approach and our boosting algorithms is essentially be due to the different loss functions employed. To apply the density-ratio trick, we first estimate the classification probabilities and then transform them into the density ratios by multiplying the sample size ratios. For implementation, we use the ada function provided in the R package ada (Culp et al., 2016). In addition to the “density-ratio trick”, we also compare our methods with two popular kernel-based density ratio estimators, namely KLEIP (Sugiyama et al., 2008) and uLSIF (Kanamori et al., 2009), implemented with the R package densratio (Makiyama, 2019).
For boosting, we assess both the FS and the GB algorithms. The learning rate $\nu$ is set to 0.01, and the maximum resolution of each tree is set to 4. The optimal number of trees is selected with the 5-fold cross-validation with the maximum number allowed set to 1000. For a fair comparison, AdaBoost is implemented under the same tuning parameters. For the generalized Bayesian additive tree model, we set the size of the ensemble to 200, following the standard implementation of the BART (Chipman et al., 2010). The MCMC sampler is initialized with trees containing only root nodes, and the size of the burn-in and the posterior sample is set to 2000 and 1000, respectively. In both types of the proposed models, we introduce 31 equally spaced cut points in each dimension. Further details on the implementation are provided in Supplementary Materials C.
For each scenario, we evaluate the performance under two sample size settings: (i) $n_{0}=n_{1}=5000$ (the “balanced” setting) and (ii) $n_{0}=9000$, $n_{1}=1000$ (the “unbalanced” setting). As a measure of accuracy, we use the following symmetrized squared mean error for the true density ratio $r^{*}$ and the estimated ratio $\hat{r}$:
$\frac{1}{2n_{0}}\sum^{n_{0}}_{i=1}\left(\log r^{*}(x_{i}^{(0)})-\log\hat{r}(x_{i}^{(0)})\right)^{2}+\frac{1}{2n_{1}}\sum^{n_{1}}_{i=1}\left(\log r^{*}(x_{i}^{(1)})-\log\hat{r}(x_{i}^{(1)})\right)^{2},$
and for each setting we calculate the average MSE based on 50 simulated data sets.
@@
We simulate data sets using the following three scenarios which represent different types of distributional differences between $X_{0}$ and $X_{1}$: